OpenStudy (anonymous):

What is the eleventh term in the sequence 256, 128, 64, 32, . . . ? please show me the steps

OpenStudy (jamesj):

Hmm, let's see now 256/128 = 2 and 128/64 = 2 and 64/32 = 2 Is there a pattern here?

OpenStudy (anonymous):

sorry i wasn't here for a while

OpenStudy (jamesj):

ok, do you see the pattern?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it geometric sequence

OpenStudy (jamesj):

yes. So what's the next term?

OpenStudy (anonymous):

16

OpenStudy (jamesj):

Right. So if x_j is the jth term \[ x_j = ar^{j-1} \] what are the values of a and r?

OpenStudy (jamesj):

x_1 = a = 256 ...

OpenStudy (anonymous):

thanks . i got it

OpenStudy (jamesj):

so what's the final answer?

OpenStudy (anonymous):

is it 1/4

OpenStudy (jamesj):

yep

OpenStudy (anonymous):

thanks